OpenStudy (anonymous):

Solve the equation (if possible) and check your solution. ((4x-3)/6)+(x/4)=(x-2)

OpenStudy (anonymous):

where are you stuck?

hero (hero):

\[\frac{4x - 3}{6} + \frac{x}{4} = x - 2\] Multiply both sides by 24 to get \[4(4x - 3) + 6x = 24(x - 2)\]

OpenStudy (anonymous):

Thank you so much!! :)

hero (hero):

yw